# PMG: Parameterized Motion Generator for Human-like Locomotion Control

### v1

@@ I INTRODUCTION
For experimental validation, we deployed the PMG algorithm on our self-developed humanoid robot, ZERITH Z1. It is worth noting that in the current research landscape, many successful deployments have been demonstrated on the Unitree G1, largely enabled by Unitree’s outstanding efforts in system identification and calibration, which have resulted in a remarkably small sim-to-real gap. However, such commercial products remain a black box to researchers due to their proprietary internal optimizations. Since ZERITH Z1 is a prototype platform, we introduce an optimization-based sim-to-real parameter identification method for motor dynamic and static parameters, significantly reducing discrepancies between simulation and reality. As a result, PMG demonstrates, for the first time within a single system, human-like natural motion, precise responsiveness to high-dimensional control commands (velocity, orientation and height) and VR operation, as well as efficient and verifiable sim-to-real transfer. In addition to its validated performance on ZERITH Z1, our PMG framework will be made publicly available to support reproducibility and future studies.
@@
Figure 1: Overview of the Parameterized Motion Generator (PMG). The framework comprises three stages: (a) Motion Preprocessing — raw mocap data(light bars) are retargeted to the robot and segmented into parameterized motion clips(darker bars). The database contains two types of clips: dynamic clips for locomotion, which are parameterized by phase $\phi\in[0,1)$, and static clips for posture adjustment, which are parameterized by their admissible command ranges. (b) Motion Generation — high-level commands are decomposed into a dynamic (velocity) command $\mathbf{u}_{d}$ and a static (posture) command $\mathbf{u}_{s}$. The dynamic and static modules generate respective reference trajectories, which are subsequently refined by a Ground-Aware Command Optimization module. (c) RL Training / Deployment — PMG outputs are incorporated into the observation and provided to the actor and critic; the actor’s outputs are executed by low-level PD controllers and fed back to the environment.
@@ III-A1 Motion Preprocessing
    where $q$ denotes joint angles, $\dot{q}$ joint angular velocities, $v$ base linear velocity, $\omega$ base angular velocity, and $c=\{\mu,\sigma\}$ foot contact information, include contact center and range.
@@ III-A2 Parameterized Gait Generation
$w^{(x)}=\frac{|u_{v,x}|}{\sum_{i\in\{x,y,\omega\}}|u_{v,i}|+\varepsilon}$ (6)
@@ III-A3 Ground-aware Command Optimization
As discussed in Section 2, the reference motions synthesized via weighted template interpolation do not always match the input commands. For instance, in certain situations, the commanded base height and pitch may conflict, or the maximum linear and angular velocities cannot be realized simultaneously due to gait constraints. To address this, we adopt a ground-aware optimization strategy that enforces consistency between the generated motion and the commanded inputs.
@@
For each foot in contact ($c_{f}(\phi)=1$), we enforce the static constraint:
@@
This approach effectively eliminates base drift and foot suspension artifacts caused by naive template interpolation, ensuring that the motion appears reasonable. However, since only kinematic (static) constraints are considered, dynamic feasibility is not guaranteed. Therefore, reinforcement learning or a dynamics-aware optimization is still required to ensure stability during dynamic execution.
@@ III-B Network Training
The reward design follows standard motion-tracking frameworks and consists of two main components: a task reward and an energy penalty. The task reward is designed to encourage the robot to follow the commanded velocities and postures, while also tracking reference joint positions, joint velocities, and foot-ground contact patterns, while the energy penalty discourages excessive torque or energy consumption.
@@ III-C1 Motor dynamic performance calibration
We observed substantial discrepancies between real actuators and the idealized motor model used in simulation, which degrade policy transfer. To correct this at the actuator level we perform single-joint response calibration: while holding other joints fixed, we apply a diverse excitation trajectory $q_{\mathrm{cmd}}(t)$ (mixture of steps and smooth segments) to a single motor and record the real response $q_{\mathrm{real}}(t)$ for $t=1,\dots,T$. In simulation we reproduce the same excitation and denote the simulated response by $q_{\mathrm{sim}}(t;\eta)$, where the simulator parameter vector is chosen as $\eta=[K_{p},\;K_{d},\;I]^{\top}$ (proportional and derivative gains and equivalent rotor inertia; additional terms such as damping or friction may be appended if required). We seek $\eta$ that minimizes the trajectory mismatch; a practical alignment loss is:
@@ IV-B1 Experimental Configurations
Single Gait Imitation (Baseline): This configuration directly imitates a single reference walking gait, without leveraging the Parameterized Motion Generator (PMG) introduced in this work. It serves as an upper-bound baseline for performance under pure imitation learning.
@@ IV-B2 Evaluation Procedure
All the three policies are evaluated within the Isaac Gym simulator under an identical sequence of preset task commands. During execution, no external intervention is applied even in cases of severe gait deformation or falling. We employ the following quantitative metrics for evaluation:
@@ IV-B3 Results and Discussion
Policies Motion Imitation Error Sim to Real Error
@@
Physical Informed Gains 0.0747 0.0656
@@ IV-C1 Experimental Configurations
Physical Informed Gains: PD parameters are configured based on motor rotor inertia and other physical properties, following the computation formulas reported in [15], aiming to make joint responses more consistent with real hardware dynamics.
Dynamic Calibration: The hardware uses the default PD parameters, while applying motor dynamic performance calibration as described in the this work.
Dynamic+Zero-point Calibration (Ours): The hardware maintains default PD parameters.
@@ IV-D3 Results and Discussion
As shown in the Figure 3 and Table III, the robot is able to successfully complete most teleoperation tasks, achieving an overall high success rate. However, due to the unavailability of dexterous hands and the absence of haptic feedback, failures still occur in certain cases (e.g., collisions or loss of balance).
@@ V CONCLUSIONS
In this work, we propose a Parametrized Motion Generator (PMG) for humanoid robots and validate it on our in-house platform, ZERITH Z1, through a systematic sim-to-real pipeline. Experiments are conducted on both locomotion and teleoperation tasks. The results demonstrate that PMG can generate natural and human-like reference motions, exhibiting smooth gait transitions and large-range whole-body posture adjustments.
Nevertheless, PMG, as a primarily data-driven approach, PMG remains heavily dependent on high-quality reference data and is thus more suited for periodic, walking-related behaviors, while its ability to handle complex, non-periodic motions such as dancing is still restricted. To address these challenges, future work will focus on incorporating richer dynamics constraints, improving the integration of imitation learning and RL, and expanding the diversity of reference datasets to enhance generalization and robustness.

### v2

@@ I INTRODUCTION
For experimental validation, we deployed the PMG algorithm on our self-developed humanoid robot, ZERITH Z1. It is worth noting that in the current research landscape, many successful deployments have been demonstrated on the Unitree G1, largely enabled by Unitree’s outstanding efforts in system identification and calibration, which have resulted in a remarkably small sim-to-real gap. However, such commercial products remain largely opaque to researchers due to proprietary internal optimizations. Since ZERITH Z1 is a prototype platform, we introduce an optimization-based sim-to-real parameter identification method for motor dynamic and static parameters, significantly reducing discrepancies between simulation and reality. As a result, PMG demonstrates, for the first time within a single system, human-like natural motion, precise responsiveness to high-dimensional control commands (velocity, orientation, and height) and VR operation, as well as efficient and verifiable sim-to-real transfer. In addition to its validated performance on ZERITH Z1, our PMG framework will be made publicly available to support reproducibility and future studies.
@@
Figure 1: Overview of the Parameterized Motion Generator (PMG). The framework comprises three stages: (a) Motion Preprocessing — raw mocap data (light bars) are retargeted to the robot and segmented into parameterized motion clips (darker bars). The database contains two types of clips: dynamic clips for locomotion, which are parameterized by phase $\phi$, and static clips for posture adjustment, which are parameterized by their admissible command ranges. (b) Motion Generation — high-level commands are decomposed into a dynamic (velocity) command $\mathbf{u}_{d}$ and a static (posture) command $\mathbf{u}_{s}$. The dynamic and static modules generate respective reference trajectories, which are subsequently refined by a Ground-Aware Command Optimization module. (c) RL Training / Deployment — PMG outputs are incorporated into the observation and provided to the actor and critic; the actor’s outputs are executed by low-level PD controllers and fed back to the environment.
@@ III-A1 Motion Preprocessing
    where $q$ denotes joint angles, $\dot{q}$ joint angular velocities, $v$ base linear velocity, $\omega$ base angular velocity, and $c=\{\mu,\sigma\}$ denotes foot contact information, including contact center and contact range.
@@ III-A2 Parameterized Gait Generation
$w^{(x)}=\frac{|u_{d,x}|}{\sum_{i\in\{x,y,\omega\}}|u_{d,i}|+\varepsilon}$ (6)
@@ III-A3 Ground-aware Command Optimization
As discussed in Section III-A2, the reference motions synthesized via weighted template interpolation do not always match the input commands. For instance, in certain situations, the commanded base height and pitch may conflict, or the maximum linear and angular velocities cannot be realized simultaneously due to gait constraints. To address this, we adopt a ground-aware optimization strategy that enforces consistency between the generated motion and the commanded inputs.
@@
For each foot in contact ($c_{f}(\phi)=1$), we enforce the following static constraint:
@@
This approach effectively eliminates base drift and foot suspension artifacts caused by naive template interpolation, ensuring kinematic consistency with ground-contact constraints. However, since only kinematic (static) constraints are considered, dynamic feasibility is not guaranteed. Therefore, reinforcement learning or a dynamics-aware optimization is still required to ensure stability during dynamic execution.
@@ III-B Network Training
The reward design follows standard motion-tracking frameworks and consists of two main components: a task reward and an energy penalty. The task reward is designed to encourage the robot to follow the commanded velocities and postures, while also tracking reference joint positions, joint velocities, and foot-ground contact patterns. The energy penalty discourages excessive torque or energy consumption.
@@ III-C1 Motor dynamic performance calibration
We observed substantial discrepancies between real actuators and the idealized motor model used in simulation, which degrade policy transfer. To correct this at the actuator level, we perform single-joint response calibration: while holding other joints fixed, we apply a diverse excitation trajectory $q_{\mathrm{cmd}}(t)$ (mixture of steps and smooth segments) to a single motor and record the real response $q_{\mathrm{real}}(t)$ for $t=1,\dots,T$. In simulation we reproduce the same excitation and denote the simulated response by $q_{\mathrm{sim}}(t;\eta)$, where the simulator parameter vector is chosen as $\eta=[K_{p},\;K_{d},\;I]^{\top}$ (proportional and derivative gains and equivalent rotor inertia; additional terms such as damping or friction may be appended if required). We seek $\eta$ that minimizes the trajectory mismatch; a practical alignment loss is:
@@ IV-B1 Experimental Configurations
Single Gait Imitation Policy: This configuration directly imitates a single reference walking gait, without leveraging the Parameterized Motion Generator (PMG) introduced in this work. It serves as an upper-bound baseline for performance under pure imitation learning.
@@ IV-B2 Evaluation Procedure
All three policies are evaluated within the Isaac Gym simulator under an identical sequence of preset task commands. During execution, no external intervention is applied even in cases of severe gait deformation or falling. We employ the following quantitative metrics for evaluation:
@@ IV-B3 Results and Discussion
Policies Motion Imitation Error Sim-to-Real Error
@@
Physically Informed Gains 0.0747 0.0656
@@ IV-C1 Experimental Configurations
Physically Informed Gains: PD parameters are configured based on motor rotor inertia and other physical properties, following the computation formulas reported in [15], aiming to make joint responses more consistent with real hardware dynamics.
Dynamic Calibration: The hardware uses the default PD parameters, while applying motor dynamic performance calibration as described in this work.
Dynamic+Zero-point Calibration (Ours): Using both motor dynamic calibration and motor zero-point calibration.
@@ IV-D3 Results and Discussion
As shown in Figure 3 and Table III, the robot is able to successfully complete most teleoperation tasks, achieving an overall high success rate. However, due to the unavailability of dexterous hands and the absence of haptic feedback, failures still occur in certain cases (e.g., collisions or loss of balance).
@@ V CONCLUSIONS
In this work, we propose a Parameterized Motion Generator (PMG) for humanoid robots and validate it on our in-house platform, ZERITH Z1, through a systematic sim-to-real pipeline. Experiments are conducted on both locomotion and teleoperation tasks. The results demonstrate that PMG can generate natural and human-like reference motions, exhibiting smooth gait transitions and large-range whole-body posture variations.
Nevertheless, as a primarily data-driven approach, PMG remains heavily dependent on high-quality reference data and is thus more suited for periodic, walking-related behaviors, while its ability to handle complex, non-periodic motions such as dancing is still restricted. To address these challenges, future work will focus on incorporating richer dynamics constraints, improving the integration of imitation learning and RL, and expanding the diversity of reference datasets to enhance generalization and robustness.